# Ensemble-Based Unsupervised Discontinuous Constituency Parsing by Tree Averaging

### v1

@@ Abstract
We address unsupervised discontinuous constituency parsing, where we observe a high variance in the performance of the only previous model. We propose to build an ensemble of different runs of the existing discontinuous parser by averaging the predicted trees, to stabilize and boost performance. To begin with, we provide comprehensive computational complexity analysis (in terms of P and NP-complete) for tree averaging under different setups of binarity and continuity. We then develop an efficient exact algorithm to tackle the task, which runs in a reasonable time for all samples in our experiments. Results on three datasets show our method outperforms all baselines in all metrics; we also provide in-depth analyses of our approach.
@@ 1 Introduction
The constituency tree, a hierarchy of words and phrases shown in Figure 1, is an important parse structure (Carnie, 2007). Researchers have proposed various approaches to address unsupervised constituency parsing, such as latent-variable methods (Clark, 2001; Petrov and Klein, 2007; Kim et al., 2019a) and rule-based systems (Cao et al., 2020; Li and Lu, 2023). Shayegh et al. (2023) reveal that different unsupervised parsers have low correlations with each other. They further propose an ensemble approach based on dynamic programming to boost performance.
@@
In this work, we propose an ensemble method to discontinuous constituency parsing by tree averaging, inspired by Shayegh et al. (2023), who address the ensemble of continuous constituency parsing by CYK-like dynamic programming (Younger, 1967). However, the previous approach is not directly applicable to our scenario because of the discontinuous structures in our setup. Consequently, the seemingly small change in the setting (continuous vs. discontinuous) leads to a completely different landscape of problems, requiring advanced algorithmic analysis and development.
@@ 2.1 Unsupervised Discontinuous Constituency Parsing
$R_{1}:S(x)\rightarrow A(x)$ $A\in\mathcal{N}_{1}$
$R_{2}:A(xy)\rightarrow U(x)U^{\prime}(y)$ $A\in\mathcal{N}_{1}$
$R_{3}:A(xyz)\rightarrow U(y)B(x,z)$ $A\in\mathcal{N}_{1}$
$R_{4}:A(x,y)\rightarrow U(x)U^{\prime}(y)$ $A\in\mathcal{N}_{2}$
$R_{5}:A(xy,z)\rightarrow U(x)B(y,z)$ $A\in\mathcal{N}_{2}$
$R_{6}:A(xy,z)\rightarrow U(y)B(x,z)$ $A\in\mathcal{N}_{2}$
$R_{7}:A(x,yz)\rightarrow U(y)B(x,z)$ $A\in\mathcal{N}_{2}$
$R_{8}:A(x,yz)\rightarrow U(z)B(x,y)$ $A\in\mathcal{N}_{2}$
$R_{9}:T(w)\rightarrow w$ $T\in\mathcal{P},w\in\Sigma$
@@
In our work, we observe that TN-LCFRS exhibits high variance in performance, especially for discontinuous constituents (Figure 2). In addition, there is a negative correlation between continuous and discontinuous performance among different runs: some are better at continuous constituents, while others are better at discontinuous ones.
Recently, Shayegh et al. (2023) show that an ensemble model can utilize different expertise of existing continuous parsers and smooth out their noise. A natural question in our scenario is: Can we build an ensemble of discontinuous parsers?
@@ 2.2 Averaging over Constituency Trees
We adopt the tree-averaging notion in Shayegh et al. (2023), which suggests the ensemble output should be the tree with the highest average $F_{1}$-score against the ensemble components (referred to as individuals):
$T^{*}=\operatorname*{argmax}_{T\in\mathcal{T}}\sum_{k=1}^{K}F_{1}(T,T_{k})$ (1)
where $\mathcal{T}$ is search space given an input sentence, and $T_{k}$ is the parse tree predicted by the $k$th individual. The $F_{1}$ score is commonly used for evaluating constituency parsers, and is also our ensemble objective, given by $F_{1}(T_{\text{prd}},T_{\text{ref}})=\frac{2|C(T_{\text{prd}})\cap C(T_{\text{ref}})|}{|C(T_{\text{prd}})|+|C(T_{\text{ref}})|}$, where $T_{\text{prd}}$ and $T_{\text{ref}}$ denote the predicted and reference trees, respectively, while $C(T)$ represents the set of constituents in a tree $T$.
It is noted that existing unsupervised parsers can only produce binary trees in both continuous and discontinuous settings (Shen et al., 2018a; Kim et al., 2019a; Yang et al., 2023). The binary property asserts that, given a length-$n$ sentence, $|C(T_{k})|=2n-1$ for every individual $k$, thus simplifying Eqn. (1) to
$T^{*}=\operatorname*{argmax}_{T\in\mathcal{T}}\frac{\sum_{c\in C(T)}{\mathpzc h}(c)}{|C(T)|+2n-1}$ (2)
Here, ${\mathpzc h}(c)$ counts the occurrences of a constituent $c$ in the trees predicted by the individuals. We call ${\mathpzc h(c)}$ the hit count of $c$.
@@
Shayegh et al. (2023) present a polynomial-time dynamic programming (DP) algorithm to average continuous binary trees, where the outputs are restricted to binary trees as well; here, continuous trees can be seen as having max fan-out $1$. For non-binary outputs, the DP table can be augmented by an additional axis whose size is bounded by the sentence length $n$. To handle discontinuity, we may augment the DP table with additional axes based on the maximum fan-out $F$. Overall, the time complexity of the DP algorithm is $\mathcal{O}(n^{4F+1})$; thus, Problem 1 is in P. See Appendix A.1 for the detailed proof. ∎
A curious question then is whether we have polynomial-time algorithms if dropping the assumptions of input being binary and having bounded fan-out. Having non-binary inputs is intriguing since the above DP relies on Eqn. (2), which only holds for binary inputs.
@@
What is the exact complexity category of averaging binary trees with unbounded fan-out?
@@
In the rest of this section, we develop a more general search algorithm that works at the level of Open Problem 1 but also solves Problem 1 more efficiently than high-order polynomial DP.
@@ 2.3 Our Search Algorithm
To solve the tree-averaging problem, we first convert it into an equivalent graph problem for clarity. Specifically, we construct an undirected graph $G$ with vertices being all possible constituents (although they can be largely pruned), weighted by corresponding hit counts, as defined after Eqn. (2). We put an edge between two vertices if and only if their corresponding constituents can coexist within a constituency tree, that is to say, they are either disjoint or inclusive of one another. Then, we formulate the below graph-search problem, which can solve the original tree-search problem for ensemble discontinuous parsing.
@@
The average constituency tree $T^{*}$ in Eqn. (2) is polynomial-time reducible from $Q^{*}$ in Problem 4 with $\alpha_{1}=0$ and $\alpha_{2}=2n-1$, if $G$ is built as above.
@@
Suppose the constituents corresponding to a clique $Q$ in the graph $G$ form a constituency tree, denoted by $T_{Q}$. We have $\sum_{v\in Q}w(v)=\sum_{c\in T_{Q}}{\mathpzc h}(c)$ by the construction of $G$. Thus, the objective function of Problem 4 is
$f(Q;0,2n-1)$ $=\frac{\sum_{v\in Q}w(v)+0}{|Q|+2n-1}$ (3)
$=\frac{\sum_{c\in C(T_{Q})}{\mathpzc h}(c)}{|C(T_{Q})|+2n-1}$ (4)
which is the same as the $\operatorname{argmax}$ objective in Eqn. (2). It is easy to see that each constituency tree corresponds to a clique in $G$. Therefore, the corresponding constituency tree to $Q^{*}$, guaranteed by Lemma 1, maximizes Eqn. (2). In other words, we have $T^{*}=T_{Q^{*}}$. ∎
@@ 2.4 Candidate Constituents Pruning
For every constituent $c\in C(T^{*})\backslash P$, where $P\subseteq C(T^{*})$, we have
${\mathpzc h}(c)>\min_{|P|\leq j\leq 2n-2}\frac{\sum_{i=1}^{j-|P|}\lambda^{+}_{i}\;+\;\sum_{c^{\prime}\in P}{\mathpzc h}(c^{\prime})}{j+2n-1}$ (5)
where $\lambda^{+}_{i}$ is the $i$th smallest positive hit count.
@@
To analyze the worst-case complexity, we notice that single words are constituents that must occur in all individuals. At least, we may set $P=\{c:c\text{ is a single word}\}$, and Theorem 5 yields a lower bound of
${\mathpzc h}(c)>\min_{n\leq j\leq 2n-2}\;\frac{0+nK}{j+2n-1}>\frac{nK}{4n}=\frac{K}{4}$ (6)
Then, we can find an upper bound for the number of candidates kept for the search process. Since the sum of all hit counts is always $(2n-1)K$, the number of candidates with a higher hit count than the threshold in Eqn. (6) is bounded by
$|V|\leq\Big{\lfloor}\frac{(2n-1)K}{K/4}\Big{\rfloor}<8n$ (7)
This shows that our pruning mechanism, profoundly, reduces the graph size from $\mathcal{O}(2^{n})$ to $\mathcal{O}(n)$, which in turn reduces the overall time complexity of our tree averaging from $\mathcal{O}(2^{2^{n-1}+2n})$ to $\mathcal{O}(2^{\frac{n}{2}}n^{2})$ in the worst-case scenario.
@@ 3.1 Settings
Baselines (four runs ${}^{\dagger}$
@@
3 N-PCFG${}^{(45)}$ $40.8_{\pm 0.5}$   – 0.00.0\ \ 0.00.0 $39.5_{\pm 0.4}$   – 0.00.0\ \ 0.00.0 $40.1_{\pm 3.9}$   – 0.00.0\ \ 0.00.0
4 C-PCFG${}^{(45)}$ $39.1_{\pm 1.9}$   – 0.00.0\ \ 0.00.0 $38.8_{\pm 1.3}$   – 0.00.0\ \ 0.00.0 $37.9_{\pm 3.4}$   – 0.00.0\ \ 0.00.0
5 TN-PCFG${}^{(4500)}$ $45.4_{\pm 0.5}$   – 0.00.0\ \ 0.00.0 $44.7_{\pm 0.6}$   – 0.00.0\ \ 0.00.0 $44.3_{\pm 6.4}$   – 0.00.0\ \ 0.00.0
6 N-LCFRS${}^{(45)}$ $33.7_{\pm 2.8}$   – $\ \ 2.0_{\pm 0.8}$ $32.7_{\pm 1.8}$   – $\ \ 1.2_{\pm 0.8}$ $36.9_{\pm 1.5}$   – $\ \ 0.9_{\pm 0.8}$
7 C-LCFRS${}^{(45)}$ $36.7_{\pm 1.5}$   – $\ \ 2.7_{\pm 1.4}$ $35.2_{\pm 1.2}$   – $\ \ 1.7_{\pm 1.1}$ $36.9_{\pm 3.7}$   – $\ \ 2.2_{\pm 1.0}$
8 TN-LCFRS${}^{(4500)}$ $46.1_{\pm 1.1}$   – $\ \ 8.0_{\pm 1.1}$ $45.4_{\pm 0.9}$   – $\ \ 6.1_{\pm 0.8}$ $45.6_{\pm 2.3}$   – $\ \ 8.9_{\pm 1.5}$
Individuals: TN-LCFRS${}^{(4500)}$
@@
Table 2: Main results. ${}^{\dagger}$uoted from Yang et al. (2023). $*{}^{*}$$p$-value $<0.05$ in an Improved Nonrandomized Sign test (Starks, 1979) against the best ensemble individual in each metric, indicated by underline. ${}^{**}$$p$-value $<0.01$.
@@ 3.2 Results and Analyses
In our study, we observe TN-LCFRS behaves inconsistently in different runs: some runs are good at continuous constituents (Row 11), while others are good at discontinuous ones (Row 12); both may disagree with the $F_{1}^{\text{overall}}$-best run (Row 10).
Our ensemble approach (Rows 14) achieves $F_{1}^{\text{cont}}$ and $F_{1}^{\text{disco}}$ scores comparable to, or higher than, all ensemble individuals. This eventually leads to a much higher $F_{1}^{\text{overall}}$ score, with a $p$-value $<0.01$ in an Improved Nonrandomized Sign (INS) test (Starks, 1979) with the binomial comparison over sentence-level performance. [^7] We chose INS test because it can properly handle a large number of ties (neutral cases), when they do not necessarily suggest the equivalency of the models but the inadequacy of test samples (e.g., no discontinuous constituents in a sentence). Results are consistent on all three datasets.
@@
Figure 4: Wall clock run time of tree-averaging algorithms on LASSY for different sentence lengths.
Figure 4: Wall clock run time of tree-averaging algorithms on LASSY for different sentence lengths.
Figure 5: Effectiveness of pruning on LASSY for (a) different sentence lengths, and (b) different numbers of ensemble individuals. Note that the orange dashed line does not fit the range of $y$-axis in (b).

### v2

# Tree-Averaging Algorithms for Ensemble-Based Unsupervised Discontinuous Constituency Parsing

@@ Abstract
We address unsupervised discontinuous constituency parsing, where we observe a high variance in the performance of the only previous model in the literature. We propose to build an ensemble of different runs of the existing discontinuous parser by averaging the predicted trees, to stabilize and boost performance. To begin with, we provide comprehensive computational complexity analysis (in terms of P and NP-complete) for tree averaging under different setups of binarity and continuity. We then develop an efficient exact algorithm to tackle the task, which runs in a reasonable time for all samples in our experiments. Results on three datasets show our method outperforms all baselines in all metrics; we also provide in-depth analyses of our approach. [^1] Code available at https://github.com/MANGA-UOFA/ TAA4EUDCP
[…]
scaledalign {adjustwidth}-.05-.05
[…]
scaledalignnoadjust
[…]
Tree-Averaging Algorithms for
Ensemble-Based Unsupervised Discontinuous Constituency Parsing
Behzad Shayegh   Yuqiao Wen   Lili Mou Dept. Computing Science & Alberta Machine Intelligence Institute, University of Alberta Canada CIFAR AI Chair the.shayegh@gmail.comyq.when@gmail.comdoublepower.mou@gmail.com
@@ 1 Introduction
A constituency tree, a hierarchy of words and phrases shown in Figure 1, is an important parse structure (Carnie, 2007). Researchers have proposed various approaches to address unsupervised constituency parsing, such as latent-variable methods (Clark, 2001; Petrov and Klein, 2007; Kim et al., 2019a) and rule-based systems (Cao et al., 2020; Li and Lu, 2023). In our previous work (Shayegh et al., 2024), we reveal that different unsupervised parsers have low correlations with each other, and further propose an ensemble approach based on dynamic programming to boost performance.
@@
In this work, we propose an ensemble method to discontinuous constituency parsing by tree averaging, following our previous work (Shayegh et al., 2024), in which we address the ensemble of continuous constituency parsing by CYK-like dynamic programming (Younger, 1967). However, the previous approach is not directly applicable to this paper because of the discontinuous structures in our setup. Consequently, the seemingly small change in the setting (continuous vs. discontinuous) leads to a completely different landscape of problems, requiring advanced algorithmic analysis and development.
@@
To sum up, our main contributions conclude: (1) proposing an ensemble approach to unsupervised discontinuous constituency parsing, (2) theoretically analyzing the computational complexity of various tree-averaging settings, and (3) conducting experiments on benchmark datasets to verify the effectiveness of our approach.
@@ 2.1 Unsupervised Discontinuous Constituency Parsing
In our work, we observe that TN-LCFRS exhibits high variance in performance, especially for discontinuous constituents (Figure 2). In addition, there is a negative correlation between performance on continuous and discontinuous constituents among different runs: some are better at continuous constituents, while others are better at discontinuous ones.
In our previous work (Shayegh et al., 2024), we show that an ensemble model can utilize different expertise of existing continuous parsers and smooth out their noise. A natural question now is: Can we build an ensemble of discontinuous constituency parsers?
@@ 2.2 Averaging over Constituency Trees
We adopt the tree-averaging notion in Shayegh et al. (2024), which suggests the ensemble output should be the tree with the highest average $F_{1}$-score against the ensemble components (referred to as individuals): {scaledalign} T^* = argmax_T ∈T ∑_k=1^K F_1(T, T_k) where $\mathcal{T}$ is the search space given an input sentence, and $T_{k}$ is the parse tree predicted by the $k$th individual. The $F_{1}$ score is commonly used for evaluating constituency parsers, and is also our ensemble objective, given by $F_{1}(T_{\text{prd}},T_{\text{ref}})=\frac{2|C(T_{\text{prd}})\cap C(T_{\text{ref}})|}{|C(T_{\text{prd}})|+|C(T_{\text{ref}})|}$, where $T_{\text{prd}}$ and $T_{\text{ref}}$ denote the predicted and reference trees, respectively, while $C(T)$ represents the set of constituents in a tree $T$.
It is noted that existing unsupervised parsers can only produce binary trees in both continuous and discontinuous settings (Shen et al., 2018a; Kim et al., 2019a; Yang et al., 2023). The binary property asserts that, given a length-$n$ sentence, $|C(T_{k})|=2n-1$ for every individual $k$, thus simplifying Eqn. (2.2) to {scaledalign} T^* = argmax_T ∈Tc ∈C(T)h(c)|C(T)|+2n-1 Here, ${\mathpzc h}(c)$ counts the occurrences of a constituent $c$ in the trees predicted by the individuals. We call ${\mathpzc h(c)}$ the hit count of $c$.
@@
In our previous work (Shayegh et al., 2024), we present a polynomial-time dynamic programming (DP) algorithm to average continuous binary trees, where the outputs are restricted to binary trees as well; here, continuous trees can be seen as having max fan-out $1$. For non-binary outputs, the DP table can be augmented by an additional axis whose size is bounded by the sentence length $n$. To handle discontinuity, we may augment the DP table with additional axes based on the maximum fan-out $F$. Overall, the time complexity of the DP algorithm is $\mathcal{O}(n^{4F+1})$; thus, Problem 1 is in P. See Appendix A.1 for the detailed proof. [^4] Corro (2023) develops a CYK-like algorithm for span-based nested named-entity recognition that takes discontinuous and non-binary structures into account. However, their algorithm limits its search space to constituents that contain at most one smaller multi-word constituent inside, making it not applicable to our scenario.
A follow-up question then is whether there is a polynomial-time algorithm when we relax the assumptions of input being binary and having bounded fan-out. Having non-binary inputs is intriguing since the above DP relies on Eqn. (2.2), which only holds for binary inputs.
@@
What is the exact complexity category (P or NP-complete) of averaging binary trees with unbounded fan-out?
@@
In the rest of this section, we develop a more general search algorithm that works at the level of Open Problem 1 but also solves Problem 1 more efficiently in practice than high-order polynomial DP.
@@ 2.3 Our Search Algorithm
To solve the tree-averaging problem, we first convert it into an equivalent graph problem for clarity. Specifically, we construct an undirected graph $G$ with vertices being all possible constituents (although they can be largely pruned), weighted by corresponding hit counts, as defined after Eqn. (2.2). We put an edge between two vertices if and only if their corresponding constituents can coexist within a constituency tree, that is to say, they are either disjoint or inclusive of one another. Then, we formulate the below graph-search problem, which can solve the original tree-search problem for ensemble discontinuous parsing.
@@
The average constituency tree $T^{*}$ in Eqn. (2.2) is polynomial-time reducible from $Q^{*}$ in Problem 4 with $\alpha_{1}=0$ and $\alpha_{2}=2n-1$, if $G$ is built as above.
@@
Suppose the constituents corresponding to a clique $Q$ in the graph $G$ form a constituency tree, denoted by $T_{Q}$. We have $\sum_{v\in Q}w(v)=\sum_{c\in T_{Q}}{\mathpzc h}(c)$ by the construction of $G$. Thus, the objective function of Problem 4 is {scaledalign} f(Q; 0, 2n-1) &= ∑v∈Qw(v) + 0|Q|+2n-1
= ∑c ∈C(TQ)h(c)|C(TQ)|+2n-1 which is the same as the $\operatorname{argmax}$ objective in Eqn. (2.2). It is easy to see that each constituency tree corresponds to a clique in $G$. Therefore, the corresponding constituency tree to $Q^{*}$, guaranteed by Lemma 1, maximizes Eqn. (2.2). In other words, we have $T^{*}=T_{Q^{*}}$. ∎
@@ 2.4 Candidate Constituents Pruning
For every constituent $c\in C(T^{*})\backslash P$, where $P\subseteq C(T^{*})$, we have {scaledalign} h(c) > min_|P| j 2n-2 i=1j-|P|λ+i+  c’ Ph(c’)j+2n-1 where $\lambda^{+}_{i}$ is the $i$th smallest positive hit count.
@@
To analyze the worst-case complexity, we notice that single words are constituents that must occur in all individuals. At least, we may set $P=\{c:c\text{ is a single word}\}$, and Theorem 5 yields a lower bound of {scaledalign} h(c) > min_n ≤j ≤2n-2  0 + nKj+2n-1 > nK4n = K4 Then, we can find an upper bound for the number of candidates kept for the search process. Since the sum of all hit counts is always $(2n-1)K$, the number of candidates with a higher hit count than the threshold in Eqn. (6) is bounded by {scaledalign} |V| ≤(2n-1)KK/4 < 8n This shows that our pruning mechanism, profoundly, reduces the graph size from $\mathcal{O}(2^{n})$ to $\mathcal{O}(n)$, which in turn reduces the overall time complexity of our tree averaging from $\mathcal{O}(2^{2^{n-1}+2n})$ to $\mathcal{O}(2^{\frac{n}{2}}n^{2})$ in the worst-case scenario.
@@ 3.1 Settings
Baselines (four runs each)
@@
3 N-PCFG(45) $40.8_{\pm 0.5}$     – 0.00.0\ \ 0.00.0 $39.5_{\pm 0.4}$     – 0.00.0\ \ 0.00.0 $40.1_{\pm 3.9}$     – 0.00.0\ \ 0.00.0
4 C-PCFG(45) $39.1_{\pm 1.9}$     – 0.00.0\ \ 0.00.0 $38.8_{\pm 1.3}$     – 0.00.0\ \ 0.00.0 $37.9_{\pm 3.4}$     – 0.00.0\ \ 0.00.0
5 TN-PCFG(4500) $45.4_{\pm 0.5}$     – 0.00.0\ \ 0.00.0 $44.7_{\pm 0.6}$     – 0.00.0\ \ 0.00.0 $44.3_{\pm 6.4}$     – 0.00.0\ \ 0.00.0
6 N-LCFRS(45) $33.7_{\pm 2.8}$     – $\ \ 2.0_{\pm 0.8}$ $32.7_{\pm 1.8}$     – $\ \ 1.2_{\pm 0.8}$ $36.9_{\pm 1.5}$     – $\ \ 0.9_{\pm 0.8}$
7 C-LCFRS(45) $36.7_{\pm 1.5}$     – $\ \ 2.7_{\pm 1.4}$ $35.2_{\pm 1.2}$     – $\ \ 1.7_{\pm 1.1}$ $36.9_{\pm 3.7}$     – $\ \ 2.2_{\pm 1.0}$
8 TN-LCFRS(4500) $46.1_{\pm 1.1}$     – $\ \ 8.0_{\pm 1.1}$ $45.4_{\pm 0.9}$     – $\ \ 6.1_{\pm 0.8}$ $45.6_{\pm 2.3}$     – $\ \ 8.9_{\pm 1.5}$
Individuals: TN-LCFRS(4500)
@@
Table 2: Main results. Quoted from Yang et al. (2023). $p$-value $<0.05$ in an Improved Nonrandomized Sign test (Starks, 1979) against the best ensemble individual in each metric, indicated by underline. ∗∗$p$-value $<0.01$.
@@ 3.2 Results and Analyses
In our study, we observe TN-LCFRS behaves inconsistently in different runs: some runs are good at continuous constituents (Row 11), while others are good at discontinuous ones (Row 12); both may disagree with the best run according to $F_{1}^{\text{overall}}$ (Row 10).
Our ensemble approach (Rows 14) achieves $F_{1}^{\text{cont}}$ and $F_{1}^{\text{disco}}$ scores comparable to, or higher than, all individuals. This eventually leads to a much higher $F_{1}^{\text{overall}}$ score, with a $p$-value $<0.01$ in an Improved Nonrandomized Sign (INS) test (Starks, 1979) with the binomial comparison over sentence-level performance. [^9] We chose INS test because it can properly handle a large number of ties (neutral cases), when they do not necessarily suggest the equivalency of the models but the inadequacy of test samples (e.g., no discontinuous constituents in a sentence). Results are consistent on all three datasets.
@@
Figure 4: Effectiveness of pruning on LASSY for (a) different sentence lengths, and (b) different numbers of ensemble individuals. Note that the dashed orange line does not fit the range of $y$-axis in (b).
Figure 5: Wall clock run time of tree-averaging algorithms on LASSY for different sentence lengths, using an Intel(R) Core(TM) i9-9940X (@3.30GHz) CPU.